Ticket Number: HEA-NOC/20041004-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: TCD
Ticket Opened: 20041004 07:25 UTC+1 Problem Start: 20041003 21:05 UTC+1
Ticket Update: 20041007 10:05 UTC+1 Problem End:
Hits on TCD primary access, Colt GE LE-001892.
Loss of connectivity with TCD.
20041004 07:19 MN Check BH access router, logs show outage from 21:05 to
23:30, and now from 06:34. Report to Colt, ticket #448068 issued.
20041004 07:54 MN Line comes up at 07:48. Check site router: it shows
cycle at 07:48. Advise Colt.
20041004 11:11 BN Line has been up since 07:48, line appears to be running
clean. Crashdump info has been passed on to Lancomms and Cisco TAC are
investigating. Connection should still be considered at risk.
20041006 09:57 BN Router instability occured this morning during
maintenance window. Connection was eventually restored and crash
information is being gathered.
20041006 12:40 BN Cisco have confirmed that the router requires
replacement hardware. This will take place either today or tomorrow.
20041006 15:13 BN This work will now take place at 08:00 tomorrow morning.
The router will be powered down and the equipment replaced. A window of
one hour is being given for this work, however the expected outage to TCD
is c. 15 minutes. No traffic will flow to or from TCD during this
20041007 09:51 BN TCD site router has now had hardware replaced and is
back in production, however the IOS upgrade once again caused issues and
so after a longer than expected outage the system is still running its
original IOS. The case has been referred back to Cisco TAC. Traffic to
TCD is flowing normally.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20041004-1